# Multi-modal Misinformation Detection: Approaches, Challenges and Opportunities

### v4

@@ 3. Multi-modal model study
Extracted features and the way they are fused play an important role in the model architecture. In fact, the model-based and feature-based studies are closely related through fusion strategies which makes the demarcation of these two studies very difficult. Hence, in this section, first we discuss common fusion strategies as the point of connection between the two studies. Furthermore, we categorize existing works based on the machine learning techniques that are exploited by each work. Specifically, we classify them into two main categories: 1) classic machine learning, and 2) deep learning-based solutions. In this section, we discuss each category in detail. .
@@ 3.2. Classic machine learning solutions
For instance, a work by Shu et al. (K. Shu and Liu, 2019), proposes individual embedded representations for text, user-user interactions, user-article interactions and publisher-article interactions, and defines a joint optimization problem leveraging these individual representations. Finally, Alternative Least Square (ALS) algorithm is utilized to solve the proposed optimization problem.
In another work, Abdali et al. propose HiJoD (Abdali et al., 2020), which encodes three different aspects of an article i.e., article text, context of social sharing behaviors and host website/domain features into individual embeddings and extracts shared structures of these embeddings by canceling out the unshared structures using a principled tensor-based framework. The extracted shared structures are utilized for article classification. In this work, the classification performance of the algebraic joint model i.e., HiJoD is compared with the naive concatenation of embedding representations and it is shown that tensor-based representation is more effective in capturing the nuance patterns of the joint structure.
More recently, Meel et al. (Hakak et al., 2021), have proposed an ensemble framework which leverages text embedding, a score calculated by cosine similarity between image caption and news body, and noisy images. Despite the fact that some of the modules of this model e.g., text embedding generator leverage deep attention-based architecture, the classification process is done via a classic ensemble technique i.e. max voting.
@@ 3.3.6. Foundation models and prompt-based techniques
Paper Concat. Attention Generative GNN Cross-modal Cue Prompting
Concatenation-based
 (Singhal et al., 2019)
 (Shu et al., 2019a)
 (Alonso-Bartolome and Segura-Bedmar, 2021)
 (Qi et al., 2021)
 (Rezayi et al., 2021)
 (Raj and Meel, 2022)
 (Jin et al., 2017)
 (Lu and Li, 2020)
 (Qian et al., 2021)
Attention-based
 (Messina et al., 2021)
 (Sachan et al., 2021)
 (Kumari and Ekbal, 2021)
 (Jing et al., 2021)
 (Wang et al., 2021c)
 (Song et al., 2021a)
 (Wang et al., 2022)
 (Jain et al., 2022)
 (Wang et al., 2018)
Generative Networks
 (Khattar et al., 2019)
 (Zeng et al., 2020)
 (Jaiswal et al., 2021)
 (Silva et al., 2021b)
 (Zhou et al., 2022)
 (Wang et al., 2021a)
GNN
 (Wang et al., 2020)
 (Dhawan et al., 2022)
Cross-modal Cue Aware
 (Zhou et al., 2020b)
 (Giachanou et al., 2020)
 (Xue et al., 2021)
 (Shang et al., 2021)
 (Singhal et al., 2021)
 (Biamby et al., 2021)
 (Choi and Ko, 2022)
Prompt-based
 (Cao et al., 2023b)
 (Cui et al., 2023)
 (Cao et al., 2023a)
 (Jiang et al., 2023)
@@ 4. Multi-modal data study
A summary and a side by side comparison of the aforementioned datasets are illustrated in Fig. 4 and Table 2. As demonstrated, most of these datasets are small in size, are annotated with binary labels, extracted from limited resources such as Twitter, and only contain limited modalities i.e., text and image.
Figure 4. Number of news articles by dataset.
@@
Fakeddit (Nakamura et al., 2020) 1,063,106 2,3,6 image,text Reddit
682,996 samples are multi-modal
NewsBag (Jindal et al., 2020) 215,000 2 image, text
TTrain: Wall Street&Onion.Test: TheRealNews & ThePoke
This dataset is highly imbalanced. There are only 15,000 fake samples.
NewsBag++ (Jindal et al., 2020) 589,000 2 image,text
Train: Wall Street&Onion. Test: TheRealNews & ThePoke
Same as NewsBag but fake samples are synthetic samples created by augmentation techniques
MM-COVID (Li et al., 2020a) 11,173 2
image,text,social context
Twitter
3,981 fake samples and 7,192 real samples
ReCOVery (Zhou et al., 2020a) 2,029 2 text,image Twitter
Imbalanced with ratio of 2:1 real vs.fake
CoAID (Cui and Lee, 2020) 5,216 2 image,text Twitter
Consists of 296,752 user engagements (926 social platforms)
MMCoVaR (Chen et al., 2021)
2,593 articles & 24,184 tweets
2
image,text,social context
Twitter
Tweets as labeled as reliable,inconclusive and unreliable
N24News (Wang et al., 2021b) 60,000 24 image,text New York Times
All samples are real from 24 different categories
MuMiN (Nielsen and McConville, 2022) 10,920 3 image,text Twitter
Consists of 10,920 articles and 6,573 images.

### v6

@@ 1. Introduction
We conducted our literature search across multiple databases, including IEEE Xplore, ACM Digital Library. and Google Scholar, using a combination of keywords related to our research focus. The inclusion criteria for the papers were defined by relevance to the research question, publication date within the last ten years to ensure timeliness, and peer-reviewed status to guarantee quality. The selection process involved an initial screening of titles and abstracts, followed by a full-text review to confirm that each paper met our stringent criteria. This methodical approach ensures that the included papers provide a diverse yet focused perspective on the subject, offering readers a succinct and informative summary of current knowledge in the field. We acknowledge the importance of transparency in our literature selection process and outline these steps to clarify the criteria and rationale behind our choices.
@@ 3. Multi-modal model study
Extracted features and the way they are fused play an important role in the model architecture. In fact, the model-based and feature-based studies are closely related through fusion strategies which makes the demarcation of these two studies very difficult. Hence, in this section, first we discuss common fusion strategies as the point of connection between the two studies. Furthermore, we categorize existing works based on the machine learning techniques that are exploited by each work. Specifically, we classify them into two main categories: 1) classic machine learning, and 2) deep learning-based solutions. In this section, we discuss each category in detail.
@@ 3.2. Classic machine learning solutions
For instance, a work by Shu et al. (K. Shu and Liu, 2019), proposes individual embedded representations for text, user-user interactions, user-article interactions and publisher-article interactions, and defines a joint optimization problem leveraging these individual representations. Finally, they apply “Non-convex Optimization” solution via the Alternating Least Squares (ALS) algorithm to solve the proposed optimization problem.
In another work, Abdali et al. (Abdali et al., 2020) propose an “Algebraic Joint Structure” algorithm called HiJoD, which encodes three different aspects of an article: the article text, the context of social sharing behaviors, and host website/domain features. These aspects are transformed into individual embeddings, and shared structures among these embeddings are extracted using a principled tensor-based framework. By canceling out the unshared structures, the extracted shared structures are then utilized for article classification. The classification performance of the algebraic joint model, HiJoD, is compared with the “Naive Embeddings Concatenation” of embedding representations. The results demonstrate that the tensor-based representation is more effective in capturing the nuanced patterns of the joint structure.
More recently, Meel et al. (Hakak et al., 2021), have proposed an “Ensemble Framework” which leverages text embedding, a score calculated by cosine similarity between image caption and news body, and noisy images. Despite the fact that some of the modules of this model e.g., text embedding generator leverage deep attention-based architecture, the classification process is done via a classic ensemble technique i.e. max “Voting”.
@@ 3.3.6. Foundation models and prompt-based techniques
Paper Concat. Attention Generative GNN Cross-modal Cue Prompting Primary Focus
(Singhal et al., 2019) Concatenation
(Shu et al., 2019a) Concatenation
(Alonso-Bartolome and Segura-Bedmar, 2021) Concatenation
(Qi et al., 2021) Concatenation
(Rezayi et al., 2021) Concatenation
(Raj and Meel, 2022) Concatenation
(Jin et al., 2017) Attention Mech.
(Lu and Li, 2020) Attention Mech.
(Qian et al., 2021) Attention Mech.
(Messina et al., 2021) Attention Mech.
(Sachan et al., 2021) Attention Mech.
(Kumari and Ekbal, 2021) Attention Mech.
(Jing et al., 2021) Attention Mech.
(Wang et al., 2021c) Attention Mech.
(Song et al., 2021a) Attention Mech.
(Wang et al., 2022) Attention Mech.
(Jain et al., 2022) Attention Mech.
(Wang et al., 2018) Generative Net.
(Khattar et al., 2019) Generative Net.
(Zeng et al., 2020) Generative Net.
(Jaiswal et al., 2021) Generative Net.
(Silva et al., 2021b) Generative Net.
(Zhou et al., 2022) Generative Net.
(Wang et al., 2021a) Generative Net.
(Wang et al., 2020) GNN
(Dhawan et al., 2022) GNN
(Zhou et al., 2020b) Cross-Modal Cue
(Giachanou et al., 2020) Cross-Modal Cue
(Xue et al., 2021) Cross-Modal Cue
(Shang et al., 2021) Cross-Modal Cue
(Singhal et al., 2021) Cross-Modal Cue
(Biamby et al., 2021) Cross-Modal Cue
(Choi and Ko, 2022) Cross-Modal Cue
(Cao et al., 2023b) Prompting
(Cui et al., 2023) Prompting
(Cao et al., 2023a) Prompting
(Jiang et al., 2023) Prompting
Multimodal Misinformation Detection Classic Machine Learning Deep Learning Embeddings Concatenation Algebraic Joint Structure Ensemble Learning & Voting Non-convex Optimization Concatenation-based Networks Attention-based Networks Graph Neural Networks Cross-modal Discordance Aware Networks Foundation Models & Prompt-based Techniques
Figure 3. An overview of the multi-modal solutions.
@@ 4. Multi-modal data study
Table 2. A comprehensive evaluation of classification performance on the Fakeddit dataset focused on accuracy. The study explored different image-text embedders. (Nakamura et al., 2020)
2-way 3-way 6-way
Type Text Image Validation Test Validation Test Validation Test
Text+Image InferSent VGG16 0.8655 0.8658 0.8618 0.8624 0.8130 0.8130
InferSent EfficientNet 0.8328 0.8339 0.8259 0.8256 0.7266 0.7280
InferSent ResNet50 0.8888 0.8891 0.8855 0.8863 0.8546 0.8526
BERT VGG16 0.8694 0.8699 0.8644 0.8655 0.8177 0.8208
BERT EfficientNet 0.8334 0.8318 0.8265 0.8255 0.7258 0.7272
BERT ResNet50 0.8929 0.8909 0.8905 0.8900 0.8600 0.8588
@@
A summary and a side by side comparison of the aforementioned datasets are illustrated in Fig. 4 and Table 3. As demonstrated, most of these datasets are small in size, are annotated with binary labels, extracted from limited resources such as Twitter, and only contain limited modalities i.e., text and image. Additionally, Table 2 showcases a comparison and evaluation of various methods’ performance on the Fakeddit dataset (Nakamura et al., 2020) [^31] The table is based on the work in (Nakamura et al., 2020) . .
@@
Fakeddit (Nakamura et al., 2020) 1,063,106 2,3,6 image,text Reddit 682,996 samples are multi-modal
NewsBag (Jindal et al., 2020) 215,000 2 image, text TTrain: Wall Street&Onion.Test: TheRealNews & ThePoke This dataset is highly imbalanced. There are only 15,000 fake samples.
NewsBag++ (Jindal et al., 2020) 589,000 2 image,text Train: Wall Street&Onion. Test: TheRealNews & ThePoke Same as NewsBag but fake samples are synthetic samples created by augmentation techniques
MM-COVID (Li et al., 2020a) 11,173 2 image,text,social context Twitter 3,981 fake samples and 7,192 real samples
ReCOVery (Zhou et al., 2020a) 2,029 2 text,image Twitter Imbalanced with ratio of 2:1 real vs.fake
CoAID (Cui and Lee, 2020) 5,216 2 image,text Twitter Consists of 296,752 user engagements (926 social platforms)
MMCoVaR (Chen et al., 2021) 2,593 articles & 24,184 tweets 2 image,text,social context Twitter Tweets as labeled as reliable,inconclusive and unreliable
N24News (Wang et al., 2021b) 60,000 24 image,text New York Times All samples are real from 24 different categories
MuMiN (Nielsen and McConville, 2022) 10,920 3 image,text Twitter Consists of 10,920 articles and 6,573 images.